### v1

# Comparing Hallucination Detection Metrics for Multilingual Generation

Abstract
While many automatic hallucination detection techniques have been proposed for English texts, their effectiveness in multilingual contexts remains unexplored. This paper aims to bridge the gap in understanding how these hallucination detection metrics perform on non-English languages. We evaluate the efficacy of various detection metrics, including lexical metrics like ROUGE and Named Entity Overlap and Natural Language Inference (NLI)-based metrics, at detecting hallucinations in biographical summaries in many languages; we also evaluate how correlated these different metrics are to gauge whether they measure the same phenomena. Our empirical analysis reveals that while lexical metrics show limited effectiveness, NLI-based metrics perform well in high-resource languages at the sentence level. In contrast, NLI-based metrics often fail to detect atomic fact hallucinations. Our findings highlight existing gaps in multilingual hallucination detection and motivate future research to develop more robust detection methods for LLM hallucination in other languages.
1 Introduction
Large Language Models (LLMs) have brought about remarkable advances in text generation. However, they are still prone to factuality hallucination - the generated text conflicts with established world knowledge Huang et al. (2023); Zhang et al. (2023). Despite considerable research efforts towards understanding and detecting hallucinations, the focus has predominantly been on English texts Huang et al. (2023); Zhang et al. (2023); Ji et al. (2023). This emphasis has resulted in a significant knowledge gap regarding hallucinations in multilingual contexts, and it is currently unclear whether the methods developed for detecting and addressing hallucinations in English are effective or even applicable in multilingual settings.
In this paper, we address the highlighted issues by evaluating the effectiveness of various metrics (initially proposed for English generation) within a multilingual context. Our approach involves a comparative analysis of traditional lexical metrics, such as ROUGE Lin (2004) and Named Entity Overlap, alongside Natural Language Inference (NLI) metrics. Additionally, we compare reference-based metrics with pairwise metrics that are based on the consistency among generated samples; we also present a correlation study between these automated metrics and human factuality verification. These evaluations shed light on the efficacy of these metrics in non-English languages and aim to establish a more robust framework for hallucination detection in multilingual language models.
We empirically evaluate different hallucination detection techniques in the multilingual context and find that:
  • Lexical overlap metrics (e.g., ROUGE, named entity overlap) do not correlate with NLI-based metrics in detecting factual hallucinations in both reference-free and reference-based settings.
  • Pairwise NLI-based metrics strongly correlate with reference-based metrics in high-resource languages, but this significantly diminishes in low-resource languages.
  • Automatic NLI-based metrics effectively identify sentence-level hallucinations in high-resource languages when compared to human evaluations. However, their performance diminishes when assessing simpler atomic facts.
These findings indicate several open problems for multilingual hallucination detection. While traditional lexical overlap methods and pairwise comparisons of multiple generated texts are more accessible approaches for low-resource languages, they are often inadequate for hallucination detection. This highlights that the effectiveness of hallucination detection is closely tied to the availability and quality of language resources, following the trend observed in English that detection accuracy depends on natural language understanding abilities Manakul et al. (2023); Min et al. (2023). Our work points to a substantial gap in hallucination detection in multilingual contexts, necessitating focused future research to bridge this divide.
2 Analyzing Methods for Multilingual Hallucination Detection
Hallucination refers to machine-generated content that is not faithful to a reference. This reference could be source text, preceding generated context, or world knowledge Ji et al. (2023); Zhang et al. (2023). Hallucinations can be further classified into two types: verifiable hallucinations, which happen when the generated content directly contradicts the reference, and unverifiable hallucinations, meanwhile, refer to generated content that cannot be verified with the source material.
In this paper, we measure the efficacy of different metrics on detecting multilingual hallucinations. We focus on biography generation, a domain that is particularly sensitive to factual accuracy and coherence Min et al. (2023); Dhuliawala et al. (2023). We test a suite of automatic metrics, each of which caters to a different aspect of factual generation: ROUGE Lin (2004), named entity overlap, and Natural Language Inference (NLI)-based methods.
2.1 Multilingual Biography Generation
Table 1: Generation quality statistics for BLOOMZ-mt. The languages in bold are the most frequent in the ROOTS pretraining corpus Laurençon et al. (2022), and the underlined languages are covered in the xP3mt fine-tuning dataset Muennighoff et al. (2023). ”FLang.” refers to the most frequently generated language for each prompt language.
Lang. #Token #Sent. Valid % FLang. Acc.
en 78.3 2.64 99.97 en 96.0
zh 115.8 4.30 100.00 zh 92.43
es 62.8 2.01 100.00 es 92.33
fr 71.3 2.24 100.00 fr 93.23
vi 45.6 1.66 98.92 vi 71.67
id 46.3 1.76 98.30 en 36.45
de 63.3 2.33 99.58 en 2.79
it 58.1 1.94 99.76 en 3.31
ja 50.3 1.97 90.73 zh 21.85
bg 17.4 1.15 86.74 en 13.69
ro 9.6 0.93 80.24 en 2.68
sv 7.6 0.51 40.73 en 1.79
th 14.8 0.81 77.08 th 94.96
ru 10.2 0.68 55.49 ru 50.44
uk 5.7 0.40 35.24 uk 41.87
fa 3.2 0.13 10.80 ur 29.90
fi 1.7 0.11 9.76 fi 34.52
ko 2.0 0.09 8.37 ko 47.30
hu 0.8 0.05 6.28 pt 14.36
Avg. 34.8 1.35 31.65 - 50.01
Inspired by prior work measuring factuality in English Min et al. (2023), we generate parallel biographies in different languages. The generated texts are then compared against a reference text (for reference-based metrics) and other generated samples (pairwise metrics) to detect hallucinations.
This section characterizes the generation quality of these biographies (Table 1). We consider the average length of each biography (in tokens and sentences), along with estimates of how accurate the generation language is to the prompt language, as in some cases, multilingual LMs will generate continuations in an unexpected language Kang et al. (2023); Bawden and Yvon (2023).
The length of the generated texts varies notably across languages. High-resource languages like English and French average 78.3 and 115.8 tokens per response, respectively. However, mid-resource languages such as Thai tend to generate much shorter biographies and incomplete sentences, and low-resource languages fare even worse (for instance, Ukrainian averages just 5.7 tokens and 0.40 sentences), demonstrating the significant gap in generation abilities across languages.
We assess the accuracy of the generated languages through three metrics: the percentage of valid generations that is detectable for the langdetect package (Valid %) [^1] Some generations are incomplete, or their language is undetectable., the most frequently generated language for a given target language (Flang), and the accuracy of generated language out of the valid generations (Acc.). For high-resource languages like English, Chinese, Spanish, and French, the models generally generate text in the correct language; however, for the languages highlighted with an underwave the model generates in the wrong language the majority of the time. Often, this is due to the model generating in a closely related high-resource language. For languages such as Italian and Bulgarian, many inaccurate generations are in English. Similarly, Japanese generations often switch to Chinese when mistakes occur. Languages with more distinctive linguistic features —such as Thai’s unique script—facilitate more accurate model generations.
2.2 Automatic Methods for Multilingual Factuality Detection
Following the language verification and ruling out examples where the generations are in the wrong languages, the next phase involves the detection of hallucinations in a multilingual environment by assessing the consistency between a target generation and either a reference text or its other generations. This section considers various automatic metrics for detecting hallucinations in long-form generations.
ROUGE
The ROUGE metric is employed to assess the token-level similarity between texts. We consider the ROUGE 1 (R1), 2 (R2), L (RL), and Lsum (RLsum) scores of the generated text against the reference.
Named Entity Overlap (NEO)
We hypothesize that the sets of named entities in the gold and generated text will differ if there is hallucination in the generation Nan et al. (2021). We calculate the F1, precision, and recall scores of named entities between the generated and reference text as an estimate for factual hallucinations.
NLI-based Detection
Following Manakul et al. (2023) and Elaraby et al. (2023), we adopt the NLI-based zero-shot sentence-level SUMMAC ($SummaC_{zs}$) scoring system Laban et al. (2021) to evaluate hallucinations. The $SummaC_{zs}$ method was originally developed to gauge the consistency between a summary $S$ and a document $D$, by segmenting them into sentences $S_{1},\dots,S_{N}$ and $D_{1},\dots,D_{M}$ respectively. Aligning with the optimal configuration in Laban et al. (2021), we employ the $max$ operator to compute the score for a sentence. Denote $e_{S_{n}}^{D_{m}}$ and $c_{S_{n}}^{D_{m}}$ as the entailment and contradiction score for the generated sentence $S_{n}$ given the reference sentence $D_{m}$, respectively.
We define three metrics to quantify verifiable hallucination and one metric to quantify unverifiable hallucination, respectively. At sentence-level detection, for a generated sentence $S_{i}$ and a reference $D$, to detect verifiable hallucination, we define the following three metrics: $\mathbf{ENT_{S_{i}}}=\max_{m}{e_{S_{i}}^{D_{m}}}$, $\mathbf{CON_{S_{i}}}=\max_{m}{c_{S_{i}}^{D_{m}}}$, and $\mathbf{DIFF_{S_{i}}}=\max_{m}{e_{S_{i}}^{D_{m}}}-\max_{m}{c_{S_{i}}^{D_{m}}}$. To detect unverifiable hallucination, we define the following metric:
$\mathbf{UNV_{S_{i}}}=1-\max(\max_{m}{e_{S_{i}}^{D_{m}}},\max_{m}{c_{S_{i}}^{D_{m}}})$
When evaluating each of the above hallucination metrics on a generated text $\hat{t}$, we consider two settings as the reference text $t$:
Reference-based This setting compares $\hat{t}$ against the relevant biographical article in Wikipedia.
Pairwise We generate $k$ samples for each biography. In this setting, we compare $\hat{t}$ against the other generated samples for the same person and calculate the average score across all generations.
3 Experiment Setup
3.1 Dataset
Our curated dataset encompasses 19 languages: English, Spanish, Russian, Indonesian, Vietnamese, Persian, Ukrainian, Swedish, Thai, Japanese, German, Romanian, Hungarian, Bulgarian, French, Finnish, Korean, Italian, and Chinese. Using WikiData, we extract 500 human names that are covered by all of these languages on Wikipedia, based on diverse page view counts from 2022-01-01 to 2023-01-01. For our reference text, we use the Wikipedia API to obtain the full-page content. We detect instances where the LLMs generate text in an incorrect language with langdetect. [^2] APIs: https://query.wikidata.org/, https://pypi.org/project/wikipedia/, and https://pypi.org/project/langdetect/, respectively
3.2 Models and Prompting Details
In our experiments, we deploy the pretrained multilingual BLOOM model as well as the BLOOMZ-mt model, which is fine-tuned with machine-translated prompts Workshop (2023). We use nucleus decoding Holtzman et al. (2020) with $top\_p=0.9$ and generate five responses per prompt. For each evaluation language, we generate a prompt template with Google Translate. The template in English is ”Tell me a biography of ¡Name¿.”; all templates are in Appendix (Figure 2).
3.3 Detection Metric Details
ROUGE scores are calculated with TorchMetrics [^3] https://github.com/Lightning-AI/torchmetrics, and we remove all stopwords before calculating ROUGE-1. Entities are extracted with Spacy’s named entity recognizer [^4] https://spacy.io/api/entityrecognizer; we note that this tagger only covers 13 of the 19 languages considered in our experiments. For the NLI-based metric, we finetune the XLMR-large model Conneau et al. (2020) on the subset of the XNLI dataset Conneau et al. (2018) that intersects with the languages used in our experiments. The finetuned model has an average validation accuracy of 85.4% for the nine intersecting languages.
Language $\mathrm{ENT}$ $\mathrm{DIFF}$ $\mathrm{UNV}$
# examples in correct language ¿ 1,000
English 0.55 0.38 0.19
French 0.52 0.40 0.15
Chinese 0.56 0.41 0.21
Spanish 0.46 0.41 0.17
Thai 0.36 0.39 0.32
Vietnamese 0.35 0.31 0.00
Indonesian 0.28 0.31 0.09
# examples in correct language ¡ 1,000
Russian 0.16 0.21 0.11
Japanese 0.37 0.40 0.07
Ukrainian 0.23 0.19 0.17
Bulgarian 0.42 0.32 0.28
Korean 0.05 0.08 -0.01
# examples in correct language ¡ 100
Finnish 0.09 0.12 0.01
Italian 0.12 0.13 0.14
Persian 0.13 0.15 0.02
German 0.50 0.45 0.11
Romanian 0.00 0.00 0.14
Hungarian 0.24 0.21 0.11
Swedish 0.30 0.27 -0.29
Table 2: The correlation between the reference-based NLI result and the pairwise NLI result across different languages. The languages with underline are covered in the XNLI finetuing dataset. The numbers in yellow have the p-values larger than 0.05.
4 Results
In this section, we compare evaluation metrics to estimate hallucinations in our generated biographical corpus (Section 4.1). We then perform a correlation study to test whether the proposed metrics agree when hallucination occurs (Section 4.2). We find that traditional, reference-based lexical metrics (i.e., ROGUE and named entity overlap) are likely insufficient to capture hallucinations, as they rarely correlate with stronger NLI-based approaches. We also observe that pairwise metrics detect hallucinations less often in lower-resource languages when compared to reference-based metrics.
4.1 Comparison of Automatic Hallucination Metrics
Table 3: Results of different reference-based metrics for the BLOOMZ-mt model. ”-” indicates the language is not covered by the Spacy NER tool. All of the ROUGE and Named Entity Overlap (NEO) results are in percentage (%).
Language R1-F1 R1-P R1-R R2-F1 R2-P R2-R NEO-F1 NEO-P NEO-R $\mathrm{DIFF}$ $\mathrm{UNV}$
High-Resource Languages
English 1.83 87.58 0.94 0.87 47.38 0.44 4.27 53.41 2.26 -0.60 0.19
Chinese 6.43 57.34 3.76 2.07 23.22 1.17 4.69 35.27 2.79 -0.62 0.21
Spanish 2.77 85.86 1.47 1.35 49.10 0.71 3.28 48.48 1.76 -0.51 0.18
French 2.18 87.78 1.13 1.06 51.41 0.55 4.35 57.41 2.31 -0.54 0.16
Vietnamese 6.82 92.92 4.22 4.10 73.28 2.43 - - - -0.49 0.15
Indonesian 7.51 68.51 4.87 2.36 26.39 1.53 - - - -0.45 0.22
Middle-Resource Languages
German 0.38 71.34 0.19 0.11 35.22 0.05 0.83 36.06 0.42 -0.65 0.15
Italian 0.50 69.13 0.25 0.13 29.82 0.07 1.00 30.26 0.52 -0.58 0.17
Japanese 0.73 14.62 0.40 0.05 1.50 0.03 0.47 15.52 0.25 -0.72 0.21
Bulgarian 0.16 4.92 0.09 0.01 0.37 0.01 - - - -0.61 0.19
Romanian 1.02 69.75 0.53 0.42 45.79 0.22 0.39 17.47 0.20 -0.29 0.24
Swedish 0.66 86.37 0.33 0.32 77.23 0.16 1.28 45.24 0.66 -0.40 0.63
Low-Resource Languages
Thai 0.04 1.14 0.02 0.00 0.00 0.00 - - - -0.56 0.38
Russian 0.09 4.69 0.05 0.01 0.28 0.00 0.48 11.28 0.25 -0.58 0.47
Ukrainian 0.04 1.53 0.02 0.00 0.00 0.00 0.70 20.64 0.36 -0.53 0.66
Persian 0.00 0.00 0.00 0.00 0.00 0.00 - - - -0.50 0.92
Finnish 0.89 37.70 0.46 0.20 10.03 0.10 0.58 23.71 0.30 -0.59 0.91
Korean 0.18 6.58 0.09 0.01 0.88 0.00 0.24 8.48 0.12 -0.53 0.94
Hungarian 0.74 64.74 0.37 0.16 23.23 0.08 - - - -0.53 0.97
We first consider how different automatic methods for detecting hallucination perform across languages on the generated biographical data from the BLOOMZ-mt model (Table 3). High-resource languages, such as English, Chinese, Spanish, French, Vietnamese, and Indonesian, exhibit higher recall scores, which suggests that the text generated in these languages has better coverage of the corresponding Wikipedia reference content. This contrasts sharply with lower-resource languages, which demonstrate significantly diminished recall. Interestingly, languages that frequently produce incorrect language outputs (e.g., German and Italian) or often yield empty or incomplete generations (e.g., Swedish and Hungarian) still have relatively high precision scores. While these generations seem to contain few explicit hallucinations, they also exclude many facts from the reference, as indicated by their correspondingly low recall scores.
When considering the $\mathrm{DIFF}$ metrics, all languages under scrutiny resulted in negative scores, which was consistent even among the higher-resource languages like English and Chinese. This indicates a tendency towards contradictions in the generated text with their respective reference texts. For the $\mathrm{UNV}$ scores, higher and middle-resource languages (ranging from English to Romanian in the table 3) cluster within a similar scope of 0.15 to 0.25. In contrast, low-resource languages that often produce empty or incomplete generations, such as Ukrainian, Persian, Finnish, and Korean, show much higher $\mathrm{UNV}$ scores. This implies that the $\mathrm{UNV}$ metric is sensitive to incomplete text generations and missing information and may indicate the model’s generation errors beyond hallucination. As for the reference-free pairwise metrics, we observe similar trends across different languages (Table 9).
4.2 Correlation Study Across Metrics
In this section, we conduct a correlation analysis to determine how various metrics align in measuring hallucination in multilingual contexts. This includes (1) the correlation between lexical hallucination metrics and NLI-based metrics, (2) the performance of the four reference-based NLI metrics, and (3) the relationship between pairwise metrics and reference-based metrics.
Figure 1: Heat map of the Pearson Correlation between reference-based metrics averaged over high-resource languages. All the P-values are less than 0.05.
Entity Generation
Example of Annotation
Alessandro Del Piero
Gen: Alessandro Del Piero, born on September 28, 1976 in Brescia, Italy, is a former Italian professional football player who served as a forward. Wiki: Alessandro Del Piero, Italian male football player…The old Maldini, who was the head coach of the national team at the time, appointed newcomers Del Piero and Vieri as the main forwards… Comment: There are 4 facts in this sentence, with 1 contradictory hallucination and 1 unverifiable hallucination. The birth date is wrong; Wikipedia doesn’t mention the birth place; for the last entity, the evidence indirectly support it.
Example of Instruction-Conflict Hallucination
Carl Edward Sagan
Gen: Stephen Hawking (January 8, 1942 -) is a famous British physicist and cosmologist.
Example of Preceding-Context-Conflict Hallucination
Blanca of Castile
Gen: Blanca of Castile (Blanca of Castile), born in 1188 and died in 1252, was the princess of Castile, queen and regent of France. Blanca is a city in the Autonomous Community of Castilla-La Mancha, Spain.
Table 4: Examples of annotation process and types of hallucinations. Sentences highlighted in red represent verifiable hallucinations that contradict evidence found in Wikipedia; those marked in yellow denote unverifiable hallucinations which lack clear evidence for verification and content highlighted in green is considered factual.
Lexical hallucination metrics do not correlate with NLI-based metrics.
Figure 1 shows that in high-resource languages (i.e., English, Chinese, French, Spanish, Vietnamese, and Indonesian), ROUGE-1 and ROUGE-L metrics demonstrate a high degree of correlation, as do the ROUGE metrics and Named Entity Overlap (NEO). However, we generally find no correlation between lexical- and NLI-based metrics, indicating that while both lexical- and NLI-based approaches are commonly proposed as automatic methods for hallucination detection, they do not measure the same deviations from a reference text.
Reference-based NLI-based metrics.
We also observe interesting trends regarding the relationship between different NLI-based metrics (Figure 1). We find that $\mathrm{ENT}$ scores are highly correlated with the $\mathrm{DIFF}$ score, indicating that these metrics identify similar artifacts in the text. Surprisingly, no correlation is observed between $\mathrm{UNV}$ and $\mathrm{CON}$ scores. Moreover, we find a negative correlation between $\mathrm{UNV}$ and $\mathrm{CON}$ scores. This is because sentences that include verifiable hallucinations likely contradict the reference text. In contrast, sentences with information that is unsubstantiated by the reference (e.g., unverifiable) will be identified as neutral instead.
Pairwise and reference-based metrics correlate in higher-resource languages.
For high-resource languages in the XNLI finetuning dataset (English, French, Chinese, Spanish, Bulgarian, and German), we observe a higher correlation in the range of 0.35 to 0.56 for pairwise NLI metrics when it comes to detecting verifiable hallucinations by $\mathrm{ENT}$ score (Table 2). This suggests pairwise metrics can identify generated content that deviates from the reference. However, the Pearson Correlation Coefficient shows lower correlation values (in the range of 0.15 to 0.21) compared to pairwise $\mathrm{UNV}$. For lower-resource languages, such as Finnish, Italian, Persian, and others, the correlation coefficients with the entailment score are often in the range of 0.00 to 0.30 and not statistically significant, highlighting the challenge of effective hallucination detection in contexts with limited language resources.
5 Human Evaluation of Multilingual Hallucination Metrics
Currently, there are no standard multilingual hallucination detection datasets available on which to compare the considered methods. We, therefore, annotate model generations in English and Chinese; annotations are performed by native speakers of each language. They manually find all verifiable and unverifiable hallucinations by checking if the generation is supported by the Wikipedia reference at both the sentence- and atomic-fact-level. Examples of resulting annotations are shown in Table 4.
l—cccc—cccc
Sentence-Level Atomic-Fact-Level
[…]
Random - 10.84 82.86 50.00 - 52.1143.56 50.00
[…]
R2-P. 0.08 19.78 81.48 - 0.10 51.23 44.23 -
RL-P. 0.11 20.04 83.10 - 0.12 52.18 42.83 -
NEO-P. 0.14 17.49 80.84 - 0.09 53.09 45.32-
$\mathrm{DIFF}$ 0.21 38.46 89.49 68.21 0.19 57.46 54.41 56.26
$\mathrm{ENT}$ 0.31 40.32 90.86 70.97 0.23 60.71 57.48 53.47
$\mathrm{CON}$ 0.11 16.47 80.41 49.13 -0.01 51.49 52.16 50.48
Reference
R2-P. 0.21 30.05 89.08 - 0.19 53.28 46.25 -
RL-P. 0.17 28.54 85.35 - 0.13 50.31 49.93 -
NEO-P. 0.17 16.15 83.75 - 0.12 57.54 47.51 -
$\mathrm{DIFF}$ 0.34 56.11 94.14 75.00 0.31 65.85 60.90 63.19
$\mathrm{ENT}$ 0.49 65.32 94.96 78.97 0.35 68.00 63.6964.18
$\mathrm{CON}$ 0.08 31.56 87.49 69.77 -0.19 53.18 57.43 59.70
Table 5: Sentence- and atomic-fact-level verifiable hallucination detection in human evaluation. F denotes factual examples and NF denotes non-factual examples. The numbers in yellow have p-values of the correlation is larger than 0.05. Note that the correlation coefficient of $\mathrm{CON}$ should be expected as negative.

### v2

@@
-Comparing Hallucination Detection Metrics for Multilingual Generation
+Comparing Hallucination Detection Methods for Multilingual Generation
 Abstract
-While many automatic hallucination detection techniques have been proposed for English texts, their effectiveness in multilingual contexts remains unexplored.
-This paper aims to bridge the gap in understanding how these hallucination detection metrics perform on non-English languages.
-We evaluate the efficacy of various detection metrics, including lexical metrics like ROUGE and Named Entity Overlap and Natural Language Inference (NLI)-based metrics, at detecting hallucinations in biographical summaries in many languages; we also evaluate how correlated these different metrics are to gauge whether they measure the same phenomena.
-Our empirical analysis reveals that while lexical metrics show limited effectiveness, NLI-based metrics perform well in high-resource languages at the sentence level.
-In contrast, NLI-based metrics often fail to detect atomic fact hallucinations.
-Our findings highlight existing gaps in multilingual hallucination detection and motivate future research to develop more robust detection methods for LLM hallucination in other languages.
+While many hallucination detection techniques have been evaluated on English text, their effectiveness in multilingual contexts remains unknown.
+This paper assesses how well various factual hallucination detection metrics (lexical metrics like ROUGE and Named Entity Overlap, and Natural Language Inference (NLI)-based metrics) identify hallucinations in generated biographical summaries across languages.
+We compare how well automatic metrics correlate to each other and whether they agree with human judgments of factuality.
+Our analysis reveals that while the lexical metrics are ineffective, NLI-based metrics perform well, correlating with human annotations in many settings and often outperforming supervised models.
+However, NLI metrics are still limited, as they do not detect single-fact hallucinations well and fail for lower-resource languages.
+Therefore, our findings highlight the gaps in exisiting hallucination detection methods for non-English languages and motivate future research to develop more robust multilingual detection methods for LLM hallucinations.
+[^1] The code and annotated dataset will be released upon publication.
 1 Introduction
-Large Language Models (LLMs) have brought about remarkable advances in text generation.
-However, they are still prone to factuality hallucination - the generated text conflicts with established world knowledge Huang et al.
+Large Language Models (LLMs) have made remarkable advances in text generation.
+However, they are still prone to hallucinating facts, or generating text that conflicts with established world knowledge Huang et al.
 (2023); Zhang et al.
 (2023).
-Despite considerable research efforts towards understanding and detecting hallucinations, the focus has predominantly been on English texts Huang et al.
+While there has been considerable research towards detecting hallucinations in English Huang et al.
 (2023); Zhang et al.
 (2023); Ji et al.
-(2023).
-This emphasis has resulted in a significant knowledge gap regarding hallucinations in multilingual contexts, and it is currently unclear whether the methods developed for detecting and addressing hallucinations in English are effective or even applicable in multilingual settings.
-In this paper, we address the highlighted issues by evaluating the effectiveness of various metrics (initially proposed for English generation) within a multilingual context.
-Our approach involves a comparative analysis of traditional lexical metrics, such as ROUGE Lin (2004) and Named Entity Overlap, alongside Natural Language Inference (NLI) metrics.
-Additionally, we compare reference-based metrics with pairwise metrics that are based on the consistency among generated samples; we also present a correlation study between these automated metrics and human factuality verification.
-These evaluations shed light on the efficacy of these metrics in non-English languages and aim to establish a more robust framework for hallucination detection in multilingual language models.
-We empirically evaluate different hallucination detection techniques in the multilingual context and find that:
-  • Lexical overlap metrics (e.g., ROUGE, named entity overlap) do not correlate with NLI-based metrics in detecting factual hallucinations in both reference-free and reference-based settings.
-  • Pairwise NLI-based metrics strongly correlate with reference-based metrics in high-resource languages, but this significantly diminishes in low-resource languages.
-  • Automatic NLI-based metrics effectively identify sentence-level hallucinations in high-resource languages when compared to human evaluations.
-However, their performance diminishes when assessing simpler atomic facts.
-These findings indicate several open problems for multilingual hallucination detection.
-While traditional lexical overlap methods and pairwise comparisons of multiple generated texts are more accessible approaches for low-resource languages, they are often inadequate for hallucination detection.
-This highlights that the effectiveness of hallucination detection is closely tied to the availability and quality of language resources, following the trend observed in English that detection accuracy depends on natural language understanding abilities Manakul et al.
+(2023), much less focus has been given to multilingual hallucinations.
+Therefore, it is currently unclear whether the methods developed for detecting and addressing hallucinations in English are effective or even applicable in multilingual settings.
+This paper evaluates the effectiveness of various automatic metrics, initially proposed for English factual hallucination detection, within a multilingual context.
+We focus on automatic metrics requiring minimal in-language resources to perform hallucination detection, as this makes them most readily applicable to new languages; these metrics include traditional lexical metrics, such as ROUGE Lin (2004) and Named Entity Overlap, as well as Natural Language Inference (NLI) metrics.
+We also consider the differences between reference-based metrics and pairwise metrics based on the consistency among generated samples.
+To evaluate these metrics, we present correlation studies comparing these automated metrics directly, against supervised hallucination detection methods, and with human judgments of generation factuality.
+We empirically evaluate these hallucination detection techniques in the multilingual context with a new dataset of parallel biographical generations (Section 2.1).
+Our experiments find that: (1) lexical overlap metrics do not agree with NLI metrics or human judgments when detecting hallucinations in reference or pairwise settings; (2) while pairwise NLI metrics strongly correlate with reference-based ones in high-resource languages, this significantly diminishes in low-resource settings; (3) automatic NLI metrics effectively detect sentence-level hallucinations in high-resource languages when compared to human evaluations, but not when assessing atomic facts; and (4) NLI metrics outperform supervised approaches at detecting hallucinations that can be verified or refuted by the reference text, but not on unverifiable errors.
+Overall, while lexical overlap methods and pairwise comparisons of generated texts are more accessible for evaluating low-resource languages, they are often inadequate at hallucination detection.
+Additionally, while NLI-based metrics can detect factual hallucinations — and even outperform models trained on hallucination detection in some cases — these metrics perform best on high-resource languages.
+This highlights that multilingual hallucination detection performance is closely tied to the availability and quality of language resources, mirroring the trend observed in English that detection depends on natural language understanding abilities Manakul et al.
 (2023); Min et al.
 (2023).
-Our work points to a substantial gap in hallucination detection in multilingual contexts, necessitating focused future research to bridge this divide.
-2 Analyzing Methods for Multilingual Hallucination Detection
-Hallucination refers to machine-generated content that is not faithful to a reference.
-This reference could be source text, preceding generated context, or world knowledge Ji et al.
-(2023); Zhang et al.
-(2023).
-Hallucinations can be further classified into two types: verifiable hallucinations, which happen when the generated content directly contradicts the reference, and unverifiable hallucinations, meanwhile, refer to generated content that cannot be verified with the source material.
-In this paper, we measure the efficacy of different metrics on detecting multilingual hallucinations.
+This points to a substantial gap in hallucination detection in multilingual and low-resource contexts and the need for future work bridging this divide.
+2 Multilingual Hallucination Detection
+We measure the efficacy of different automatic metrics on detecting multilingual hallucinations.
 We focus on biography generation, a domain that is particularly sensitive to factual accuracy and coherence Min et al.
@@
 2.1 Multilingual Biography Generation
-Table 1: Generation quality statistics for BLOOMZ-mt.
-The languages in bold are the most frequent in the ROOTS pretraining corpus Laurençon et al.
-(2022), and the underlined languages are covered in the xP3mt fine-tuning dataset Muennighoff et al.
-(2023).
-”FLang.” refers to the most frequently generated language for each prompt language.
+Inspired by prior work measuring factuality in English Min et al.
+(2023), we generate parallel biographies in different languages.
+The generated texts are then compared against a reference text (for reference-based metrics) and other generated samples (pairwise metrics) to detect hallucinations.
+This section characterizes the generation quality of these biographies (Table 1).
+We consider the average length of each biography (in tokens and sentences), along with estimates of how accurate the generation language is to the prompt language, as in some cases, multilingual LMs will generate continuations in an unexpected language Kang et al.
+(2023); Bawden and Yvon (2023).
+The length of the generated texts varies notably across languages.
+While high-resource languages like English and French generate longer outputs, mid-resource languages such as Thai tend to generate much shorter biographies and incomplete sentences.
+Low-resource languages fare even worse (for instance, Ukrainian averages just 5.7 tokens and 0.40 sentences), demonstrating the significant gap in generation abilities across languages.
+We assess the accuracy of the generated languages through three metrics: the percentage of valid generations that is detectable for the langdetect package (Valid %), the most frequently generated language for a given target language (Flang), and the accuracy of generated language out of the valid generations (Acc.).
+For high-resource languages like English, Chinese, Spanish, and French, the models generally generate text in the correct language; however, for the languages highlighted with an underwave the model generates in the wrong language the majority of the time.
+Often, this is due to the model generating in a closely related high-resource language.
+For languages such as Italian and Bulgarian, many inaccurate generations are in English.
+Similarly, Japanese generations often switch to Chinese when mistakes occur.
+Languages with more distinctive linguistic features —such as Thai’s unique script—facilitate more accurate model generations.
 Lang. #Token #Sent. Valid % FLang. Acc.
@@
 Avg. 34.8 1.35 31.65 - 50.01
-Inspired by prior work measuring factuality in English Min et al.
-(2023), we generate parallel biographies in different languages.
-The generated texts are then compared against a reference text (for reference-based metrics) and other generated samples (pairwise metrics) to detect hallucinations.
-This section characterizes the generation quality of these biographies (Table 1).
-We consider the average length of each biography (in tokens and sentences), along with estimates of how accurate the generation language is to the prompt language, as in some cases, multilingual LMs will generate continuations in an unexpected language Kang et al.
-(2023); Bawden and Yvon (2023).
-The length of the generated texts varies notably across languages.
-High-resource languages like English and French average 78.3 and 115.8 tokens per response, respectively.
-However, mid-resource languages such as Thai tend to generate much shorter biographies and incomplete sentences, and low-resource languages fare even worse (for instance, Ukrainian averages just 5.7 tokens and 0.40 sentences), demonstrating the significant gap in generation abilities across languages.
-We assess the accuracy of the generated languages through three metrics: the percentage of valid generations that is detectable for the langdetect package (Valid %) [^1] Some generations are incomplete, or their language is undetectable., the most frequently generated language for a given target language (Flang), and the accuracy of generated language out of the valid generations (Acc.).
-For high-resource languages like English, Chinese, Spanish, and French, the models generally generate text in the correct language; however, for the languages highlighted with an underwave the model generates in the wrong language the majority of the time.
-Often, this is due to the model generating in a closely related high-resource language.
-For languages such as Italian and Bulgarian, many inaccurate generations are in English.
-Similarly, Japanese generations often switch to Chinese when mistakes occur.
-Languages with more distinctive linguistic features —such as Thai’s unique script—facilitate more accurate model generations.
-2.2 Automatic Methods for Multilingual Factuality Detection
-Following the language verification and ruling out examples where the generations are in the wrong languages, the next phase involves the detection of hallucinations in a multilingual environment by assessing the consistency between a target generation and either a reference text or its other generations.
-This section considers various automatic metrics for detecting hallucinations in long-form generations.
+Table 1: Quality statistics for BLOOMZ-mt generations.
+Languages that occur in the ROOTS pretraining corpus are in bold Laurençon et al.
+(2022), and underlined languages are in the xP3mt fine-tuning dataset Muennighoff et al.
+(2023).
+"FLang." refers to the most frequently generated language for each prompt language.
+2.2 Automatic Metrics
+After quality verification of generated samples and filtering examples where the output is in an incorrect language, we compare the efficacy of different hallucination detection metrics on the remaining generations.
+We consider automatic metrics for detecting hallucinations in long-form generations that work by assessing the consistency between a target generation and either a reference text or its other generations.
+Specifically, we focus on metrics that do not require supervised hallucination data: many languages do not have datasets available for this task, which makes these supervised methods infeasible for those settings.
+[^2] However, for completeness we include two recent supervised methods for multilingual hallucination detection in §5.
 ROUGE
-The ROUGE metric is employed to assess the token-level similarity between texts. We consider the ROUGE 1 (R1), 2 (R2), L (RL), and Lsum (RLsum) scores of the generated text against the reference.
+The ROUGE metric is employed to assess the token-level similarity between texts. We consider the generated text’s ROUGE 1 (R1) and L (RL) scores against the reference.
 Named Entity Overlap (NEO)
@@
 In this setting, we compare $\hat{t}$ against the other generated samples for the same person and calculate the average score across all generations.
+Experimental details for calculating these metrics are given in the Appendix.
 3 Experiment Setup
-3.1 Dataset
+Dataset
 Our curated dataset encompasses 19 languages: English, Spanish, Russian, Indonesian, Vietnamese, Persian, Ukrainian, Swedish, Thai, Japanese, German, Romanian, Hungarian, Bulgarian, French, Finnish, Korean, Italian, and Chinese.
-Using WikiData, we extract 500 human names that are covered by all of these languages on Wikipedia, based on diverse page view counts from 2022-01-01 to 2023-01-01.
+Using WikiData, we extract the names of 500 people who are covered by all of these languages on Wikipedia, based on diverse page view counts from 2022-01-01 to 2023-01-01.
 For our reference text, we use the Wikipedia API to obtain the full-page content.
-We detect instances where the LLMs generate text in an incorrect language with langdetect.
-[^2] APIs: https://query.wikidata.org/, https://pypi.org/project/wikipedia/, and https://pypi.org/project/langdetect/, respectively
-3.2 Models and Prompting Details
-In our experiments, we deploy the pretrained multilingual BLOOM model as well as the BLOOMZ-mt model, which is fine-tuned with machine-translated prompts Workshop (2023).
+We detect instances where the LLMs generate text in an incorrect language with langdetect, which covers all 19 languages in our experiment.
+[^3] APIs: https://query.wikidata.org/, https://pypi.org/project/wikipedia/, and https://pypi.org/project/langdetect/, respectively.
+Models and Prompting
+We generate text samples with the BLOOMZ-mt model, which is fine-tuned with machine-translated prompts Workshop (2023); at the time of our experiments, BLOOMZ-mt is the largest open-source, multilingual LM.
 We use nucleus decoding Holtzman et al.
-(2020) with $top\_p=0.9$ and generate five responses per prompt.
+(2020) with $top\_p=0.9$, which is a common and realistic configuration used in other works in LLM hallucination Liu et al.
+(2023), and generate five responses per prompt to evaluate the pairwise, intrinsic metrics.
 For each evaluation language, we generate a prompt template with Google Translate.
-The template in English is ”Tell me a biography of ¡Name¿.”; all templates are in Appendix (Figure 2).
-3.3 Detection Metric Details
-ROUGE scores are calculated with TorchMetrics [^3] https://github.com/Lightning-AI/torchmetrics, and we remove all stopwords before calculating ROUGE-1.
-Entities are extracted with Spacy’s named entity recognizer [^4] https://spacy.io/api/entityrecognizer; we note that this tagger only covers 13 of the 19 languages considered in our experiments.
-For the NLI-based metric, we finetune the XLMR-large model Conneau et al.
-(2020) on the subset of the XNLI dataset Conneau et al.
-(2018) that intersects with the languages used in our experiments.
-The finetuned model has an average validation accuracy of 85.4% for the nine intersecting languages.
+The template in English is "Tell me a biography of <Name>."; the templates translated into other languages are in Appendix (Figure 2).
+Language R1-F1 R1-P R1-R RL-F1 RL-P RL-R N-F1 N-P N-R $\mathrm{DIFF}$ $\mathrm{UNV}$ $\mathrm{ENT}$
+High-Resource Languages
+English 1.83 87.58 0.94 1.40 72.41 0.72 4.27 53.41 2.26 -0.60 0.19 0.16
+Chinese 6.43 57.34 3.76 5.59 51.73 3.26 4.69 35.27 2.79 -0.62 0.21 0.16
+Spanish 2.77 85.86 1.47 2.19 72.39 1.16 3.28 48.48 1.76 -0.51 0.18 0.28
+French 2.18 87.78 1.13 1.67 73.51 0.87 4.35 57.41 2.31 -0.54 0.16 0.25
+Vietnamese 6.82 92.92 4.22 5.34 85.87 3.21 - - - -0.49 0.15 0.34
+Indonesian 7.51 68.51 4.87 5.44 55.28 3.52 - - - -0.45 0.22 0.32
+Middle-Resource Languages
+German 0.38 71.34 0.19 0.31 67.60 0.16 0.83 36.06 0.42 -0.65 0.15 0.50
+Italian 0.50 69.13 0.25 0.42 63.77 0.21 1.00 30.26 0.52 -0.58 0.17 0.42
+Japanese 0.73 14.62 0.40 0.64 13.53 0.35 0.47 15.52 0.25 -0.72 0.21 0.26
+Bulgarian 0.16 4.92 0.09 0.15 4.91 0.08 - - - -0.61 0.19 0.50
+Romanian 1.02 69.75 0.53 1.00 69.08 0.52 0.39 17.47 0.20 -0.29 0.24 0.76
+Swedish 0.66 86.37 0.33 0.64 85.87 0.33 1.28 45.24 0.66 -0.40 0.63 0.79
+Low-Resource Languages
+Thai 0.04 1.14 0.02 0.04 1.14 0.02 - - - -0.56 0.38 0.32
+Russian 0.09 4.69 0.05 0.09 4.62 0.05 0.48 11.28 0.25 -0.58 0.47 0.40
+Ukrainian 0.04 1.53 0.02 0.03 1.52 0.02 0.70 20.64 0.36 -0.53 0.66 0.51
+Persian 0.00 0.00 0.00 0.00 0.00 0.00 - - - -0.50 0.92 0.38
+Finnish 0.89 37.70 0.46 0.80 35.61 0.41 0.58 23.71 0.30 -0.59 0.91 0.33
+Korean 0.18 6.58 0.09 0.18 6.57 0.09 0.24 8.48 0.12 -0.53 0.94 0.25
+Hungarian 0.74 64.74 0.37 62.56 23.23 0.36 - - - -0.53 0.97 0.51
+Table 2: Results of different reference-based metrics for the BLOOMZ-mt model.
+"-" indicates the language is not covered by the Spacy NER tool.
+All of the ROUGE and Named Entity Overlap (N) results are in percentage (%).
+4 Multilingual Hallucination Metrics
+This section compares how different automatic metrics estimate hallucinations in our generated biographical corpus (Section 4.1).
+We then perform a correlation study to test whether these metrics agree when hallucination occurs (Section 4.2).
+4.1 Automatic Metrics
+We first consider how different referenced-based automatic methods for detecting hallucination perform across languages on the generated biographical data from the BLOOMZ-mt model (Table 2).
+[^4] We observe similar trends on pairwise metrics (Table 10).
+We find that, unsurprisingly, these measures indicate increases in hallucination on middle- and low-resource languages (e.g., lower overlap with the reference, higher $\mathrm{UNV}$ scores).
+However, the NLI-based $\mathrm{DIFF}$ scores remain relatively stable regardless of language resourcefulness.
 Language $\mathrm{ENT}$ $\mathrm{DIFF}$ $\mathrm{UNV}$
-# examples in correct language ¿ 1,000
+# examples in correct language > 1,000
 English 0.55 0.38 0.19
@@
 Indonesian 0.28 0.31 0.09
-# examples in correct language ¡ 1,000
+# examples in correct language < 1,000
 Russian 0.16 0.21 0.11
@@
 Korean 0.05 0.08 -0.01
-# examples in correct language ¡ 100
+# examples in correct language < 100
 Finnish 0.09 0.12 0.01
@@
 The languages with underline are covered in the XNLI finetuing dataset.
-The numbers in yellow have the p-values larger than 0.05.
-4 Results
-In this section, we compare evaluation metrics to estimate hallucinations in our generated biographical corpus (Section 4.1).
-We then perform a correlation study to test whether the proposed metrics agree when hallucination occurs (Section 4.2).
-We find that traditional, reference-based lexical metrics (i.e., ROGUE and named entity overlap) are likely insufficient to capture hallucinations, as they rarely correlate with stronger NLI-based approaches.
-We also observe that pairwise metrics detect hallucinations less often in lower-resource languages when compared to reference-based metrics.
-4.1 Comparison of Automatic Hallucination Metrics
-Table 3: Results of different reference-based metrics for the BLOOMZ-mt model.
-”-” indicates the language is not covered by the Spacy NER tool.
-All of the ROUGE and Named Entity Overlap (NEO) results are in percentage (%).
-Language R1-F1 R1-P R1-R R2-F1 R2-P R2-R NEO-F1 NEO-P NEO-R $\mathrm{DIFF}$ $\mathrm{UNV}$
-High-Resource Languages
-English 1.83 87.58 0.94 0.87 47.38 0.44 4.27 53.41 2.26 -0.60 0.19
-Chinese 6.43 57.34 3.76 2.07 23.22 1.17 4.69 35.27 2.79 -0.62 0.21
-Spanish 2.77 85.86 1.47 1.35 49.10 0.71 3.28 48.48 1.76 -0.51 0.18
-French 2.18 87.78 1.13 1.06 51.41 0.55 4.35 57.41 2.31 -0.54 0.16
-Vietnamese 6.82 92.92 4.22 4.10 73.28 2.43 - - - -0.49 0.15
-Indonesian 7.51 68.51 4.87 2.36 26.39 1.53 - - - -0.45 0.22
-Middle-Resource Languages
-German 0.38 71.34 0.19 0.11 35.22 0.05 0.83 36.06 0.42 -0.65 0.15
-Italian 0.50 69.13 0.25 0.13 29.82 0.07 1.00 30.26 0.52 -0.58 0.17
-Japanese 0.73 14.62 0.40 0.05 1.50 0.03 0.47 15.52 0.25 -0.72 0.21
-Bulgarian 0.16 4.92 0.09 0.01 0.37 0.01 - - - -0.61 0.19
-Romanian 1.02 69.75 0.53 0.42 45.79 0.22 0.39 17.47 0.20 -0.29 0.24
-Swedish 0.66 86.37 0.33 0.32 77.23 0.16 1.28 45.24 0.66 -0.40 0.63
-Low-Resource Languages
-Thai 0.04 1.14 0.02 0.00 0.00 0.00 - - - -0.56 0.38
-Russian 0.09 4.69 0.05 0.01 0.28 0.00 0.48 11.28 0.25 -0.58 0.47
-Ukrainian 0.04 1.53 0.02 0.00 0.00 0.00 0.70 20.64 0.36 -0.53 0.66
-Persian 0.00 0.00 0.00 0.00 0.00 0.00 - - - -0.50 0.92
-Finnish 0.89 37.70 0.46 0.20 10.03 0.10 0.58 23.71 0.30 -0.59 0.91
-Korean 0.18 6.58 0.09 0.01 0.88 0.00 0.24 8.48 0.12 -0.53 0.94
-Hungarian 0.74 64.74 0.37 0.16 23.23 0.08 - - - -0.53 0.97
-We first consider how different automatic methods for detecting hallucination perform across languages on the generated biographical data from the BLOOMZ-mt model (Table 3).
-High-resource languages, such as English, Chinese, Spanish, French, Vietnamese, and Indonesian, exhibit higher recall scores, which suggests that the text generated in these languages has better coverage of the corresponding Wikipedia reference content.
-This contrasts sharply with lower-resource languages, which demonstrate significantly diminished recall.
-Interestingly, languages that frequently produce incorrect language outputs (e.g., German and Italian) or often yield empty or incomplete generations (e.g., Swedish and Hungarian) still have relatively high precision scores.
-While these generations seem to contain few explicit hallucinations, they also exclude many facts from the reference, as indicated by their correspondingly low recall scores.
-When considering the $\mathrm{DIFF}$ metrics, all languages under scrutiny resulted in negative scores, which was consistent even among the higher-resource languages like English and Chinese.
-This indicates a tendency towards contradictions in the generated text with their respective reference texts.
-For the $\mathrm{UNV}$ scores, higher and middle-resource languages (ranging from English to Romanian in the table 3) cluster within a similar scope of 0.15 to 0.25.
-In contrast, low-resource languages that often produce empty or incomplete generations, such as Ukrainian, Persian, Finnish, and Korean, show much higher $\mathrm{UNV}$ scores.
+The numbers in gray have the p-values larger than 0.05.
+Lexical Overlap Metrics
+We also note some specific trends within this metric type.
+For example, high-resource languages (English, Chinese, Spanish, French, Vietnamese, and Indonesian) exhibit particularly high recall scores, suggesting that the text generated in these languages has better coverage of the corresponding Wikipedia reference content.
+In contrast, lower-resource languages demonstrate significantly diminished recall.
+Interestingly, languages where BLOOMZ-mt frequently produces incorrect language outputs (e.g., German and Italian) or empty or incomplete generations (e.g., Swedish and Hungarian) maintain relatively high precision scores in the higher-quality outputs we evaluate.
+While these generations seem to contain few explicit hallucinations, they often exclude many facts from the reference, as indicated by their correspondingly low recall scores.
+NLI-based Metrics
+All languages we consider obtain negative $\mathrm{DIFF}$ scores, including higher-resource languages like English and Chinese.
+This indicates a tendency towards contradictions in the generated text with their respective reference texts — as measured by the NLI classifier.
+For the $\mathrm{UNV}$ scores, higher and middle-resource languages (ranging from English to Romanian in the table 2) fall within a similar range of 0.15 to 0.25.
+In contrast, low-resource languages that often produce empty or incomplete generations, such as Ukrainian, Persian, Finnish, and Korean, obtain much higher $\mathrm{UNV}$ scores.
 This implies that the $\mathrm{UNV}$ metric is sensitive to incomplete text generations and missing information and may indicate the model’s generation errors beyond hallucination.
-As for the reference-free pairwise metrics, we observe similar trends across different languages (Table 9).
 4.2 Correlation Study Across Metrics
-In this section, we conduct a correlation analysis to determine how various metrics align in measuring hallucination in multilingual contexts.
-This includes (1) the correlation between lexical hallucination metrics and NLI-based metrics, (2) the performance of the four reference-based NLI metrics, and (3) the relationship between pairwise metrics and reference-based metrics.
+In this section, we conduct a correlation analysis to determine whether the considered metrics agree in measuring hallucination in multilingual contexts.
+This includes (1) the correlation between lexical hallucination metrics and NLI-based metrics, (2) the agreement of the four reference-based NLI metrics, and (3) the relationship between pairwise metrics and reference-based metrics.
 Figure 1: Heat map of the Pearson Correlation between reference-based metrics averaged over high-resource languages. All the P-values are less than 0.05.
@@
 Example of Annotation
-Alessandro Del Piero
-Gen: Alessandro Del Piero, born on September 28, 1976 in Brescia, Italy, is a former Italian professional football player who served as a forward.
+Alessandro Del Piero Gen: Alessandro Del Piero, born on September 28, 1976 in Brescia, Italy, is a former Italian professional football player who served as a forward.
 Wiki: Alessandro Del Piero, Italian male football player…The old Maldini, who was the head coach of the national team at the time, appointed newcomers Del Piero and Vieri as the main forwards… Comment: There are 4 facts in this sentence, with 1 contradictory hallucination and 1 unverifiable hallucination.
 The birth date is wrong; Wikipedia doesn’t mention the birth place; for the last entity, the evidence indirectly support it.
-Example of Instruction-Conflict Hallucination
-Carl Edward Sagan
-Gen: Stephen Hawking (January 8, 1942 -) is a famous British physicist and cosmologist.
-Example of Preceding-Context-Conflict Hallucination
-Blanca of Castile
-Gen: Blanca of Castile (Blanca of Castile), born in 1188 and died in 1252, was the princess of Castile, queen and regent of France.
-Blanca is a city in the Autonomous Community of Castilla-La Mancha, Spain.
-Table 4: Examples of annotation process and types of hallucinations.
-Sentences highlighted in red represent verifiable hallucinations that contradict evidence found in Wikipedia; those marked in yellow denote unverifiable hallucinations which lack clear evidence for verification and content highlighted in green is considered factual.
+Table 4: Example of human hallucination annotations.
+Red represents verifiable hallucinations contradicting evidence in the reference (Wiki), yellow denotes unverifiable hallucinations without relevant evidence, and green is supported by the reference text.
 Lexical hallucination metrics do not correlate with NLI-based metrics.
-Figure 1 shows that in high-resource languages (i.e., English, Chinese, French, Spanish, Vietnamese, and Indonesian), ROUGE-1 and ROUGE-L metrics demonstrate a high degree of correlation, as do the ROUGE metrics and Named Entity Overlap (NEO).
+Figure 1 shows that in high-resource languages (i.e., English, Chinese, French, Spanish, Vietnamese, and Indonesian), ROUGE-1 and ROUGE-L metrics demonstrate a high degree of correlation, and Named Entity Overlap (NEO) correlates reasonably well with ROGUE precision metrics.
 However, we generally find no correlation between lexical- and NLI-based metrics, indicating that while both lexical- and NLI-based approaches are commonly proposed as automatic methods for hallucination detection, they do not measure the same deviations from a reference text.
 Reference-based NLI-based metrics.
-We also observe interesting trends regarding the relationship between different NLI-based metrics (Figure 1).
-We find that $\mathrm{ENT}$ scores are highly correlated with the $\mathrm{DIFF}$ score, indicating that these metrics identify similar artifacts in the text.
-Surprisingly, no correlation is observed between $\mathrm{UNV}$ and $\mathrm{CON}$ scores.
+We also observe interesting trends regarding the relationship between different NLI-based metrics (bottom right-hand corner of Figure 1).
+We find that $\mathrm{ENT}$ scores are highly (inversely) correlated with the $\mathrm{DIFF}$ score, indicating that these metrics identify similar artifacts in the text.
 Moreover, we find a negative correlation between $\mathrm{UNV}$ and $\mathrm{CON}$ scores.
@@
 In contrast, sentences with information that is unsubstantiated by the reference (e.g., unverifiable) will be identified as neutral instead.
-Pairwise and reference-based metrics correlate in higher-resource languages.
-For high-resource languages in the XNLI finetuning dataset (English, French, Chinese, Spanish, Bulgarian, and German), we observe a higher correlation in the range of 0.35 to 0.56 for pairwise NLI metrics when it comes to detecting verifiable hallucinations by $\mathrm{ENT}$ score (Table 2).
-This suggests pairwise metrics can identify generated content that deviates from the reference.
-However, the Pearson Correlation Coefficient shows lower correlation values (in the range of 0.15 to 0.21) compared to pairwise $\mathrm{UNV}$.
-For lower-resource languages, such as Finnish, Italian, Persian, and others, the correlation coefficients with the entailment score are often in the range of 0.00 to 0.30 and not statistically significant, highlighting the challenge of effective hallucination detection in contexts with limited language resources.
-5 Human Evaluation of Multilingual Hallucination Metrics
-Currently, there are no standard multilingual hallucination detection datasets available on which to compare the considered methods.
-We, therefore, annotate model generations in English and Chinese; annotations are performed by native speakers of each language.
-They manually find all verifiable and unverifiable hallucinations by checking if the generation is supported by the Wikipedia reference at both the sentence- and atomic-fact-level.
-Examples of resulting annotations are shown in Table 4.
-l—cccc—cccc
+Pairwise and reference metrics do not correlate in low-resource languages.
+For high-resource languages in the XNLI finetuning dataset (English, French, Chinese, Spanish, Bulgarian, and German), we observe higher correlations ranging between 0.35 to 0.56 for pairwise and reference-based NLI metrics when it comes to detecting verifiable hallucinations by $\mathrm{ENT}$ score (Table 3).
+This suggests pairwise metrics can identify generated content that deviates from the reference and may be useful for detecting hallucinations when gold reference texts are not available.
+However, the Pearson Correlation Coefficient shows lower correlation values (in the range of 0.15 to 0.21) when comparing pairwise and reference-based $\mathrm{UNV}$, indicating a less effective capture of extrinsic hallucinations involving plausible yet unverifiable information.
+For lower-resource languages, such as Finnish, Italian, Persian, correlation with the entailment score is often not statistically significant.
+This implies that the effectiveness of pairwise hallucination metrics is limited to higher-resource languages, highlighting the challenge of effective hallucination detection in limited resource contexts.
+Question Instructions
+0. Atomic-Level Annotation Extract a sentence that contains only one simple fact.
+1. Evidence Extraction Copy and paste all relevant evidence.
+2. Total Facts Provide an approximate count of the total facts. Each date is counted as one fact, except for birthdates and death dates, which are counted as two separate facts.
+3. Verifiable Correct Facts Count the number of facts that can be verified as correct.
+4. Verifiable Contradictory Facts Count the number of facts that contradict verified information.
+5. Unverifiable Facts Count the number of facts that cannot be verified.
+6. Conflict with Preceding Context Indicate whether there is any conflict with the preceding context (True or False).
+7.
+Conflict with Instructions Determine whether there is any conflict with the instructions.
+Label it as False if the example provides (a) a biography of (b) the correct person.
+Otherwise, label it as True.
+Table 5: Instructions for manual hallucination annotations. Step 0 is only taken for atomic-level fact verification.
+5 Human Evaluation
+We manually annotate the model generations analyzed in the prior section; the annotations are performed on paired subsets of the English and Chinese generations by native speakers.
+Following the Attributable to Identified Sources (AIS) paradigm Rashkin et al.
+(2023) for measuring hallucination, annotators manually find all verifiable and unverifiable hallucinations by checking if the generated output is attributable to the Wikipedia reference at both the sentence- and atomic-fact-level.
+Table 4 shows example annotations.
+{NiceTabular}
+l|ccc|ccc
+Sentence Level Atomic-Fact Level
+[…]
+Random - 10.84 82.86 - 52.1143.56
+[…]
+R1-P. 0.08  19.78 81.48 0.10 51.23 44.23
+RL-P. 0.11  20.04 83.10 0.12  52.18 42.83
+NEO-P. 0.14  17.49 80.84 0.09 53.09 45.32
+$\mathrm{DIFF}$ 0.21 38.46 89.49 0.19 57.46 54.41
+$\mathrm{ENT}$ 0.31 40.32 90.86 0.23 60.71 57.48
+$\mathrm{CON}$ 0.11 16.47 80.41 -0.01 51.49 52.16
+Reference
+R1-P. 0.21 30.05 89.08 0.19 53.28 46.25
+RL-P. 0.17 28.54 85.35 0.13  50.31 49.93
+NEO-P. 0.17  16.15 83.75 0.12 57.54 47.51
+$\mathrm{DIFF}$ 0.34 56.11 94.14 0.31 65.85 60.90
+$\mathrm{ENT}$ 0.49 65.32 94.96 0.35 68.00 63.69
+$\mathrm{CON}$ 0.08 31.56 87.49 -0.19 53.18 57.43
+$\mathrm{mFact}$ 0.20 35.68 91.16 0.29 67.30 61.67
+$\mathrm{Seahorse}$0.17  13.25 75.40 -0.07  53.30 46.67
+Table 6: Comparison of sentence and atomic-fact verifiable hallucination metrics with the human support rate.
+F denotes factual examples and NF denotes non-factual examples.
+p-values of this correlation is larger than 0.05.
+5.1 Experimental Setup
+The authors (one per language) manually checked every sentence in the audited subset of generations, using the steps listed in Table 5.
+For atomic-fact-level annotation, a preprocessing step is taken to extract only sentences that contain a standalone proposition [^5] A standalone proposition is independently interpretable from the information contained in the assertion.
+Rashkin et al.
+(2023).
+Then, for both the sentence- and atomic-fact-level annotation, we annotate all relevant evidence sentences from the reference Wikipedia page and accumulate the counts for different types of propositions (Table 5).
+Table 9 details data statistics from this annotation process.
+Metrics
+We compare our automatic metrics presented in §2.2 with human annotations using correlation and classification; we specifically compare precision metrics because they are generally the strongest automatic measure in Section 4.1.
+For correlation, we investigate the relationship of the metrics with the support rate (SR; $N_{vs}/N_{t}$) for verifiable hallucination detection and with the unverified rate ($N_{nv}/N_{t}$) for unverifiable hallucination detection using their Pearson correlations.
+To consider the classification agreement of these metrics, we calculate the Precision-Recall area under the curve (AUC-PR) between the discretized human-annotated and automatic metrics.
+We convert human annotations into classification labels by labeling an example as factual only if all its facts are supported by evidence for verifiable hallucinations with the support rate (Table 5); for unverifiable hallucinations, we consider any sentence with at least one fact not supported or refuted by the reference to be unverified: $N_{nv}\geq 1$ (Table 5.3).
+We discretize the automatic NLI-based metrics into classification labels by setting their respective thresholds, with 0.5 for the entailment and contradictory scores and 0 for the difference between these two scores.
+The thresholds were selected based on the different degrees of tolerance for the proportions of unverifiable hallucinations in a sentence.
+We then perform classification using the discretized human judgments as gold labels.
+5.2 Automatic Metric Results
+NLI entailment outperforms lexical metrics on sentence-level verification.
+We observe low correlation between lexical metrics like ROUGE-1 (R1) and Named Entity Overlap (NEO) and the human-annotated support rate (SR; Table 5).
+Moreover, AUC-PR for these metrics exhibits minor improvements over the random classification baseline, particularly for non-factual examples (NF), underscoring the limit of lexical metrics for accurately detecting factual hallucinations.
+In contrast, the NLI-based metrics highly correlate with different measures of human verification for detecting verifiable hallucinations at the sentence level (Table 5).
+These results corroborate the NLI metric results in English in Manakul et al.
+(2023).
+Specifically, the $\mathrm{ENT}$ and $\mathrm{DIFF}$ scores perform well, while $\mathrm{CON}$ alone does not demonstrate the same agreement with human judgments.
+Pairwise metrics underperform reference-based metrics.
+When compared against human annotations, pairwise metrics always underperform the reference-based ones: their Pearson correlations, AUCF, AUCNF, and accuracies are all worse than the metrics using references, and this holds at both the sentence- and atomic-fact-levels (Table 5).
+However, we note that pairwise NLI-based metrics, particularly the entailment and difference scores, demonstrate significantly better performance than the lexical-based pairwise metrics, though they still underperform comparable reference metrics.
+This suggests pairwise NLI-based approaches remain the most useful approach in hallucination detection settings where references are unavailable.
+NLI metrics struggle to detect unverifiable hallucination and check atomic facts.
+Comparison against human judgments on unverifiable hallucination detection reveals that automatic metrics show only marginal improvements over random classification and low correlation with human annotations (Table 5.3).
+Furthermore, NLI metrics also encounter significant challenges in accurately verifying the factuality of simple atomic facts (Table 5).
+Both AUC-PR and accuracy demonstrate a marked decrease in the effectiveness of NLI-based metrics on atomic facts.
+This aligns with Luo et al.
+(2022), which also highlighted similar limitations in English NLI metrics for verification.
+This finding underscores the need for alternative metrics to address current NLI approaches’ limitations, especially in the critical area of atomic factuality evaluation.
+5.3 Supervised Metric Results
+We also consider two supervised metrics: mFACT Qiu et al.
+(2023) and seahorse-Q4 Clark et al.
+(2023).
+mFACT Qiu et al.
+(2023) automatically apply English hallucination detection approaches to new languages by first using faithfulness metrics to rank and annotate English examples, and then translating the most and least faithful samples into a set of target languages.
+They then train a classifier on each target language (including English and Chinese).
+The seahorse-Q4 metric Clark et al.
+(2023) similarly fine-tunes a mT5-large model Xue et al.
+(2021) on an attribution task, using attribution subset (Q4) of the SEAHORSE dataset.
+Supervised metrics underperform NLI-based metrics in verifiable hallucination detection.
+mFACT scores exhibit similar but slightly lower performance than NLI-based metrics at both the sentence and atomic fact levels (Table 5).
+The seahorse-Q4 approach shows comparable performance to mFACT in English (Appendix Table D).
+However, as seahorse-Q4 does not see Chinese data in training, it fails to detect verifiable hallucinations in this setting, leading to a lower overall score and highlighting the brittleness of supervised methods for hallucination detection.
+Supervised metrics detect unverifiable hallucinations more effectively.
+[…]
+This trend also holds at the atomic-fact level, where the supervised metrics outperform other methods in across all measures.
+However, they still exhibit lower correlation with this unverified rate than at the sentence level, indicating that detecting hallucination on simpler atomic facts remains challenging in the unverified setting.
+{NiceTabular}
+[…]
+[…]
+$\mathrm{UNV}$${}_{\textit{Sent}}$ 0.28 31.77 71.84
+$\mathrm{mFACT}$${}_{\textit{Sent}}$0.29 54.17 69.67
+$\mathrm{seahorse}$${}_{\textit{Sent}}$0.10 42.08 69.49
+[…]
+$\mathrm{UNV}$${}_{\textit{Atom}}$ 0.12 77.49 27.36
+$\mathrm{mFACT}$${}_{\textit{Atom}}$0.13 80.35 31.49
+$\mathrm{seahorse}$${}_{\textit{Atom}}$0.09 78.32 28.48
+Table 7: Agreement with human judgments on unverifiable hallucination detection, averaged for Chinese and English.
+UNV denotes unverifiable hallucination, and V denotes verifiable.
+Gray values have p-values $>$ 0.05.
+6 Related Work
+Factuality Hallucination Detection
+Detecting hallucinations is crucial for ensuring the reliability of machine-generated content.
+One line of work leverages uncertainty by analyzing the probability of the LLM’s output space Mielke et al.
+(2022); Kadavath et al.
+(2022); Varshney et al.
+(2023); other methods evaluate consistency between repeatly generated samples (including NLI-based approaches) Elaraby et al.
+(2023); Manakul et al.
+(2023), similar to our pairwise evaluation setting.
+Similarly, our reference-based evaluations follow works that use external evidence as reference texts and verify if the generation is supported by this reference Chern et al.
+(2023); Min et al.
+(2023).
+However, factuality hallucination evaluations are limited in multilingual settings, with existing works using prompting methods Ahuja et al.
+(2023) and machine translation Lai et al.
+(2023).
+Task-specific Multilingual Hallucination
+The main focus of multilingual hallucination evaluation methods so far has been on specific downstream tasks.
+mFACT evaluates faithfulness in summaries by transferring English judgments into target languages via machine translation Qiu et al.
+(2023).
+Similarly, Aharoni et al.
+(2022) leverages factual consistency models to improve faithfulness in multilingual summarization.
+In neural machine translation (NMT), Dale et al.
+(2022) detect and alleviate hallucinations by measuring generation similarity to the source text; Xu et al.
+(2023) similarly study source effects on hallucination with input perturbations.
+Other works Lee et al.
+(2019); Raunak et al.
+(2021) analyze the susceptibility of current NMT methods to generate hallucinations.
+7 Conclusion
+This study investigates the effectiveness of automatic metrics for detecting factual hallucinations in non-English generations by considering these metrics directly (Section 4) and comparing their predictions to human judgments and supervised detection approaches (Section 5).
+We document that while traditional lexical metrics struggle to detect hallucinations in multilingual settings, NLI-based metrics show promise in high-resource languages at the sentence level.
+However, their effectiveness diminishes when applied to atomic facts, and the reliability of NLI-based metrics is tied to the performance of NLI models, posing a significant hurdle in lower-resource languages.
+Therefore, our analysis highlights that detecting hallucinations effectively in a language is directly linked to the availability and quality of linguistic resources in that language.
+As a result, automatically detecting hallucinations in lower-resource languages remains a significant challenge for current NLP methods.
+Limitations
+This study focuses on text generation and hallucination in a specific setting (namely, generating biographical summaries with the BLOOM-mt model) to perform a controlled study on how different automatic metrics detect factual hallucinations.
+It remains an open question whether these findings hold in other generation settings, particularly when there is less reliance on factual knowledge (such as story generation).
+Additionally, portions of our experimental setup rely on automatic methods.
+Specifically, we use machine translation to construct the prompt templates, which may introduce noise.
+Furthermore, due to the unavailability of native speakers for other languages, our human evaluation and comparison against automated metrics is limited to Chinese and English.
+In the future, we would like to expand on these findings with human evaluations in other, lower-resourced languages to confirm how well the automatic detection methods hold up in these settings.
+Appendix A Qualitative Analysis of Challenging Cases in Annotation
+We identify four challenging categories of hallucination detection for annotators and NLI metrics (Table 8).
+  • Inferred: Implicit fact connections between generation and evidence.
+  • Subjective: Generation contains subjective content, which is challenging for both human annotators and fact-based NLI models.
+  • Nuanced Difference: There are subtle distinctions between evidence and generated text, which is often missed by surface-level text classification in NLI models.
+  • Temporal Information: Generation contains time-sensitive information, which requires models to have an understanding of temporal context.
+Type Example
+Inferred Gen: Alessandro Del Piero is a former Italian professional football player and a forward. Wiki: Maldini, who was the head coach of the national team, appointed Del Piero as the main forward.
+Subjective Gen: Frida Kahlo is widely regarded as the most influential painter of the 20th century.
+Nuanced Difference Gen: Louis Pasteur is known as the "Father of Modern Microbiology". Wiki: … has been honored as the "father of microbiology"…
+Temporal Information Gen: Michelle Bachelet is currently the President of Chile. Wiki: She served as President of Chile from 2006 to 2010 and from 2014 to 2018…
+Table 8: Categories of special case in annotation.
+Each category presents unique difficulties in determining the factuality of generated content with its evidence source.
+  • Inferred connection between generated content and evidence is one of the biggest challenges for both annotators and NLI models, since they need to infer the relationship or the factual basis that links them.
+This requires a deep understanding of context and the ability to draw inferences from potentially sparse or indirect evidence.
+  • Subjective content in generations poses a significant challenge because it introduces personal opinions, emotions, or interpretations that are inherently difficult to verify against factual evidence.
+For human annotators, this can lead to variability in judgments based on personal biases or interpretations.
+For NLI models, which are primarily designed for fact-based analysis, handling subjective content requires advanced understanding of sentiment, opinion, and cultural context, areas where current models may fall short.
+  • Nuanced difference between evidence and generated text highlight the limitations of surface-level text classification approaches in NLI models.
+Detecting nuanced differences demands a granular analysis of semantics, requiring models to understand context, synonyms, and slight variations in meaning.
+This challenge underscores the need for more sophisticated NLI models capable of deep semantic analysis and the importance of training annotators to pay attention to detail and understand the significance of minor discrepancies.
+  • Time-sensitive information introduces complexity because it requires both annotators and models to have an understanding of temporal context and the ability to evaluate statements within the correct time frame.
+This can be particularly challenging when information changes over time, requiring up-to-date knowledge and the ability to discern the relevance of temporal qualifiers in text.
+For NLI models, this underscores the need for dynamic knowledge bases and the ability to reason about time, which are areas where current models may lack proficiency.
+Overall, these challenges highlight the complexities involved in hallucination detection and the need for advanced capabilities in human annotators and NLI models.
+Appendix B Metric Calculation Details
+ROUGE scores are calculated with TorchMetrics [^6] https://github.com/Lightning-AI/torchmetrics, and we remove all stopwords before calculating ROUGE-1.
+Entities are extracted with Spacy’s named entity recognizer [^7] https://spacy.io/api/entityrecognizer; we note that this tagger only covers 13 of the 19 languages considered in our experiments.
+For the NLI-based metric, we finetune the XLMR-large model Conneau et al.
+(2020) on the subset of the XNLI dataset Conneau et al.
+(2018) that intersects with the languages used in our experiments.
+The finetuned model has an average validation accuracy of 85.4% for the nine intersecting languages.
+Appendix C Additional Results
+We show the pairwise metric results in Table 10. We observe similar trends as the reference-based metrics. Also, the average statistics of annotation result are shown in Table 9.
+Metric Sent-Level Atomic-Level
+# Examples 111 102
+# Words 46.21 10.21
+# Evidence 2.17 1.00
+# Total Facts 4.76 1.00
+Support Rate 0.35 0.29
+Contradictory Rate 0.15 0.24
+Unverified Rate 0.50 0.47
+Instruction-conflict Rate 0.03 0.07
+Context-conflict Rate 0.13 0.06
+Table 9: Average statistics of Chinese and English annotation data.
+Table 10: Results of different pairwise consistency metrics for the BLOOMZ-mt model.
+"-" indicates they have no coverage for the NER tool we use.
+All of the ROUGE and Named Entity Overlap results are in percentage (%).
+Language R1_F1 R1_P R1_R RL_F1 RL_P RL_R NEO_F1 NEO_P NEO_R $\mathrm{DIFF}$ $\mathrm{UNV}$
+English 12.03 14.21 14.71 5.91 7.10 7.37 4.27 53.41 2.26 -0.25 0.57
+Chinese 7.57 8.56 8.55 4.00 4.55 4.55 4.69 35.28 2.79 -0.29 0.57
+Spanish 12.49 14.45 15.05 6.21 7.31 7.68 3.28 48.48 1.76 -0.22 0.52
+German 4.42 10.47 4.48 1.98 5.51 2.04 0.83 36.06 0.42 -0.33 0.56
+Russian 0.15 0.19 0.13 0.00 0.00 0.00 0.48 11.28 0.25 -0.09 0.55
+Indonesian 4.74 6.23 6.52 1.59 2.13 2.27 - - - -0.16 0.68
+Vietnamese 11.60 14.36 15.91 6.26 7.91 8.91 - - - -0.32 0.56
+Persian 0.00 0.00 0.00 0.00 0.00 0.00 - - - -0.31 0.52
+Ukrainian 0.00 0.00 0.00 0.00 0.00 0.00 0.70 20.64 0.36 0.14 0.42
+Swedish 10.04 14.69 10.82 8.53 13.23 9.74 1.28 45.24 0.66 -0.20 0.54
+Thai 0.01 0.02 0.01 0.00 0.00 0.00 0.00 0.00 0.00 -0.19 0.68
+Japanese 0.47 0.75 0.58 0.08 0.13 0.09 0.47 15.52 0.25 -0.07 0.56
+Romanian 12.21 13.95 12.68 8.57 9.40 8.70 0.39 17.47 0.20 -0.30 0.44
+Hungarian 0.18 0.25 0.46 0.00 0.00 0.00 - - - -0.18 0.58
+Bulgarian 0.30 0.44 0.28 0.05 0.08 0.06 - - - -0.02 0.96
+French 12.02 14.03 14.61 6.26 7.43 7.76 4.35 57.41 2.31 -0.05 0.92
+Finnish 0.46 0.47 0.61 0.17 0.19 0.19 0.58 23.71 0.30 -0.04 0.94
+Korean 0.19 0.26 0.17 0.00 0.00 0.00 0.24 8.48 0.12 0.25 0.63
+Italian 4.79 7.85 5.08 2.71 4.15 2.77 1.00 30.26 0.52 -0.01 0.97
+{NiceTabular}
+[…]
+[…]
+$\mathrm{UNV}$${}_{\textit{Sent}}$ 0.27 32.70 70.31
+$\mathrm{mFACT}$${}_{\textit{Sent}}$
+-0.18 45.2763.94
+$\mathrm{seahorse}$${}_{\textit{Sent}}$-0.03 41.74 62.73
+[…]
+[…]
+$\mathrm{UNV}$${}_{\textit{Atom}}$ 0.09 73.49 20.19
+$\mathrm{mFACT}$${}_{\textit{Atom}}$0.17 79.03 32.01
+$\mathrm{seahorse}$${}_{\textit{Atom}}$0.12 78.24 30.93
+Table 11: Results of unverifiable hallucination detection in human evaluation in English, with the best agreement with humans indicated in bold.
+Sent denotes sentence-level detection, and Atom denotes atomic-fact-level detection.
+UNV denotes unverifiable hallucination, and V denotes verifiable.
+The number in gray have the p-values larger than 0.05.
+{NiceTabular}
+[…]
+[…]
+$\mathrm{UNV}$${}_{\textit{Sent}}$ 0.31 30.27 76.12
+$\mathrm{mFACT}$${}_{\textit{Sent}}$0.32 56.54 60.42
+$\mathrm{seahorse}$${}_{\textit{Sent}}$0.11 41.45 63.54
+[…]
+$\mathrm{UNV}$${}_{\textit{Atom}}$ 0.12 79.26 27.36
+$\mathrm{mFACT}$${}_{\textit{Atom}}$0.19 86.77 34.38
+$\mathrm{seahorse}$${}_{\textit{Atom}}$0.14 80.76 30.92
+Table 12: Results of unverifiable hallucination detection in human evaluation in Chinese, with the best agreement with humans indicated in bold.
+All the p-values of the correlation is less than 0.05.
+Sent denotes sentence-level detection, and Atom denotes atomic-fact-level detection.
+UNV denotes unverifiable hallucination, and V denotes verifiable.
+The number in gray have the p-values larger than 0.05.
+Appendix D Generation Prompt Templates
+We present the full set of prompt templates for all languages from Section 3 in Figure 2.
+Figure 2: Prompt templates of all languages used in generating biography. {} represents human names.
+{NiceTabular}
+l|ccc|ccc
 Sentence-Level Atomic-Fact-Level
 […]
-Random - 10.84 82.86 50.00 - 52.1143.56 50.00
-[…]
-R2-P. 0.08 19.78 81.48 - 0.10 51.23 44.23 -
-RL-P. 0.11 20.04 83.10 - 0.12 52.18 42.83 -
-NEO-P. 0.14 17.49 80.84 - 0.09 53.09 45.32-
-$\mathrm{DIFF}$ 0.21 38.46 89.49 68.21 0.19 57.46 54.41 56.26
-$\mathrm{ENT}$ 0.31 40.32 90.86 70.97 0.23 60.71 57.48 53.47
-$\mathrm{CON}$ 0.11 16.47 80.41 49.13 -0.01 51.49 52.16 50.48
 Reference
-R2-P. 0.21 30.05 89.08 - 0.19 53.28 46.25 -
-RL-P. 0.17 28.54 85.35 - 0.13 50.31 49.93 -
-NEO-P. 0.17 16.15 83.75 - 0.12 57.54 47.51 -
-$\mathrm{DIFF}$ 0.34 56.11 94.14 75.00 0.31 65.85 60.90 63.19
-$\mathrm{ENT}$ 0.49 65.32 94.96 78.97 0.35 68.00 63.6964.18
-$\mathrm{CON}$ 0.08 31.56 87.49 69.77 -0.19 53.18 57.43 59.70
-Table 5: Sentence- and atomic-fact-level verifiable hallucination detection in human evaluation.
+random - 18.34 82.56 - 54.28 48.31
+$\mathrm{DIFF}$ 0.28 49.11 92.41 0.29 63.93 62.49
+$\mathrm{ENT}$ 0.31 45.43 90.27 0.33 66.67 63.69
+$\mathrm{CON}$0.14 31.56 92.20 -0.21 60.18 56.90
+$\mathrm{mFact}$ 0.31 31.70 85.52 0.28 64.84 60.10
+$\mathrm{Seahorse}$ 0.08  29.28 85.80 0.09  60.94 57.44
+Table 13: Sentence- and atomic-fact-level verifiable hallucination detection in human evaluation in English comparing automatic metrics and human support rate, with the best agreement with humans indicated in bold.
 F denotes factual examples and NF denotes non-factual examples.
-The numbers in yellow have p-values of the correlation is larger than 0.05.
-Note that the correlation coefficient of $\mathrm{CON}$ should be expected as negative.
+p-values of this correlation is larger than 0.05.